# Ego-Exo4D: Understanding Skilled Human Activity from First- and Third-Person Perspectives

### v3

@@ 12.A   Expert Commentary
To collect expert commentary, we developed a web-based tool, which is open sourced as part of the Ego-Exo4D dataset and benchmark suite. Known as the Narrator, this application supports video playback for Ego-Exo4D skills demonstrations, records time-stamped verbal commentary, and allows exporting and viewing commented videos. As a web-based platform, the Narrator can be simply accessed through a browser, with minimal set-up and less restrictive system requirements compared to tools requiring local installation. These attributes made it efficient to onboard and manage our geographically distributed experts. We acknowledge the EPIC Narrator [27] as the open-sourced inspiration and source code for this initiative.
Commentary Analysis
We ask at least 2 and as many as 5 distinct experts to commentate any given video, to provide a variety of language and opinions. Altogether, we have collected 117,812 pieces of commentary. On average there are 7.5 pieces of commentary per minute of video, and typically the expert gives about 4 sentences of commentary every time they pause to react to the video. Table 5 shows the number of skills demonstrations with expert commentary per domain. Overall, we believe the commentary is a unique window into the skilled actions that (through language) surfaces many subtleties about the actions not evident to the untrained eye.
Domain # experts
Basketball 6
Music 6
Soccer 5
Dance 6
Bouldering 6
Health 5
Cooking 11
Bike Repair 7
Table 5: Number of experts providing commentary in each domain
12.B   Narrate and Act
The participant surveys (Appendix 11) capture the skill of the participants, but they do not provide first-hand information regarding the actions being completed and objects they used. We capture these using “Narrate and Act": a separate take recorded with the participant in which they narrate what they are doing and how they are doing it.
Participants were asked to complete the Narrate and Act take as if they are teaching someone else as part of an instructional video or how-to guide. In this way they would talk about what they were doing as they carried out an action. This is akin to the kind of narrations done by people creating how-to videos, though with less stylization and without any professional post-production editing. It is interesting to note that some participants did this naturally requiring little prompting. For some activities which were more intense, such as dancing or bouldering, we asked participants to instead narrate either just before or just after the action to reduce the difficulty of doing this live.
Compared to the third-person expert commentary (above), these narrations are first-person and delivered at the same time of the actions. Generally the commentary is richer in constructive feedback about the quality of the activity, whereas the narrate-and-act narrations are interesting for their simultaneous nature and first-person analysis of what the participant is doing. The behavior in this extra take is expected to differ from that of the non-narrated tasks, in that it is likely that the participant will complete the scenario more slowly than normal to concentrate on explaining what they were doing. However, the narrate and act takes provided benefits in reducing annotation noise and they can potentially be used for multimodal learning as is currently explored in the literature with how-to video narrations  [89, 104, 103, 9]. Table 4 (third column) and Figure 21 show examples.
12.C   Atomic Action Descriptions
Inspired by the “narrations" in Ego4D [47], our third language resource provides atomic action descriptions of the camera wearer and his or her environment, when relevant, as captured by the Ego-Exo4D videos. Compared to the expert commentaries and narrate-and-act narrations—which both emphasize the “why" and “how"—these atomic action descriptions are focused on the objective “what" of what is happening when, written in free-form text from the perspective of a third-party observer.
@@ 12.C   Atomic Action Descriptions
Table 6: Atomic action descriptions per domain statistics.
Annotation description
We present each take to the annotators as a collaged video consisting of the Aria egocentric view, left and right grayscale SLAM, four or five fixed-position exocentric cameras, and single-track composite audio; for a subset of videos, a helmet-mounted GoPro view is also available. Annotators were asked to provide a play-by-play description of what happens, as seen across any of the views. Potential contents include actions by the Aria camera wearer, other individuals interacting with the camera wearer, and relevant environmental events. Each narration is atomic and time-anchored: as much as possible, each narration should roughly be limited to one verb and have a single associated timestamp, roughly within a second of its occurrence in the video. For consistency across narrations, the Aria camera wearer in each take is referred to as “C” in all narrations (e.g. “C picks up a wrench.”). Other individuals are referred to by other letters (e.g. “Man X kicks the soccer ball back to C.”); these letter labels are not necessarily consistent across takes, but refer to the same individual within a take. Many videos are narrated by two independent human annotators, and we make each narration set separately available. Table 4 (second column) and Figure 21 show examples. See Table 6 for atomic action descriptions summary statistics.
Visibility
Because of the multi-view nature of the Ego-Exo4D capture rig, certain actions or events may not be visible across all camera feeds. While we hope Ego-Exo4D leads to increased attention toward multi-view learning, many existing systems fundamentally assume a single view at a time; if a camera does not have a view of the narrated action or event, this may lead to a confusing learning signal, or pose an impossible ask for a model to infer. Thus, we also ask that the annotators answer two additional question per narration: 1) an indicator of whether the narration is visible from the egocentric camera, and 2) which (if any) of the static exocentric cameras provide the best view. If there are multiple equally good views, annotators are free to pick any. In particular, we found this best exocentric view helpful for other Ego-Exo4D annotation efforts: the narration visibility tags played a role in exocentric view selection for both the correspondences benchmark (Section 13.A.1) and expert commentary (above), and frame selection for hand and body pose (Section 13.D).
Relation to other Ego-Exo4D annotations
While bearing some similarities to our other annotation workstreams, these atomic action descriptions exhibit some key differences:
  • Keysteps (Section 13.B.1): While some atomic descriptions are similar to certain keystep names, narrations are free-form, without intentionally imposing any shared vocabulary. Atomic descriptions also aim to capture all actions and events, regardless of relevance as immediate steps towards the scenario’s goal, and cover all domains, not just procedural.
  • Expert Commentary (Section 12.A): As atomic description annotators do not necessarily have the years of experience of the domain experts, the resulting text is often written from a layperson’s perspective, and focus on what is happening as opposed to providing critiques, analysis, or explanation. Atomic descriptions also tend to be shorter and denser than expert commentary.
  • Narrate and Act (Section 12.B): While also a language resource, narrate and act’s descriptions are expressed from a first person perspective and can share intrinsic motivation behind actions. In contrast, atomic descriptions are written from the perspective of an outside observer.
12.D   Comparison of the Language Statistics
Figure 25 shows word clouds per scenario and annotation type highlighting the differences in vocabulary and word frequency.
Contributions Statement
13   Benchmarks: Annotations and Baselines
Benchmark Annotation Type EgoExo4D v1 EgoExo4D v2
Num Takes Annotations Num Takes Annotations
Relations Manual 1028 3419 objects 1335 5566 objects
426K ego masks 742K ego masks
611K exo masks 1.1M exo masks
Keystep recognition Manual 1088 17 activities, 664 keysteps 1088 17 activities, 664 keysteps
27.6K ego segments (87h) 27.6K ego segments (87h)
143K ego+exo segments (454h) 143K ego+exo segments (454h)
Procedure understanding Manual 628 6 activities, 186 keysteps 628 6 activities, 186 keysteps
8.6K segments (30h) 8.6K segments (30h)
Proficiency estimation Semi-automatic 2987 2987 proficiency scores 2987 2987 proficiency scores
(demonstrator) (demonstrator)
Manual 912 19K “good” segments 912 19K “good” segments
20K “tips” segments 20K “tips” segments
(demonstration) (demonstration)
Ego pose (Body) Automatic 610 1.4M 3D / 5.7M 2D 2559 9.2M 3D / 46.87M 2D
Manual 511 200K 3D / 1.03M 2D 1358 376K 3D / 2M 2D
Ego pose (Hand) Automatic 749 2.1M 3D / 10.5M 2D 976 4.3M 3D / 21M 2D
Manual 352 51K 3D / 285K 2D 458 68K 3D / 340K 2D
In this section we provide details on all the benchmark task definitions, their annotations, and our baseline models and results. We cover ego-exo relation (Sec. 13.A), followed by ego-exo recognition (Sec. 13.B), proficiency estimation (Sec. 13.C), and ego body and hand pose (Sec. 13.D).
Important: to ensure fair comparisons in any future work using Ego-Exo4D, researchers need to account for 1) the precise task input-output definitions and 2) the train/test/val splits available with v1 or v2 of the annotations. Specifically, for each task, when formally defining the inputs and outputs, we also explicitly specify which inputs are excluded from use, if any. Furthermore, there are two publicly released versions of Ego-Exo4D annotations. In this paper, only v1 is used to train the baselines, and results are reported using the v1-trained models on both the v1 and v2 test sets. The v2 test set is the basis for the leaderboard now, and the larger v2 train/test sets will be used for future challenge leaderboards. Table 7 provides summary annotation statistics for all tasks of Ego-Exo4D. Again, these two points are important information for any future research done with Ego-Exo4D to ensure consistency of results in the literature to come.
13.A   Ego-Exo Relation
The family of ego-exo relation tasks deals with relating the video content across the extreme ego-exo viewpoint changes, in the form of either object-level matching (correspondence) or generation of one view from the other (translation).
13.A.1 Ego-exo correspondence
@@
We annotate pairs of temporally synchronized egocentric and exocentric videos with segmentation masks for selected object instances from six scenarios: Cooking, Bike Repair, Health, Music, Basketball and Soccer. We exclude Bouldering and Dance from this benchmark as they have limited diversity of objects. We focus on objects used by the camera-wearer at any point during the execution of the activity and that are visible in both views for at least some frames of the sequence. These masks allow us to define object-level correspondence between the views.
@@
  • Stage 1: Egocentric video annotation. Annotator watches the egocentric video and is also shown (a) text and (b) a bounding box for one of the objects annotated in the previous stage. Annotator then marks a segmentation mask for that object in all the video frames where the object is visible. Segment Anything [70] is leveraged to generate segmentation masks efficiently using only point clicks.
@@
What are the objects of interest? We focus on objects that are active at some point during the execution of the activity. These objects are not only interesting because they are essential to the activity, but they are also challenging to track, since they are moving/changing state. In particular, our annotation guidelines requested annotators to list (a) objects that the camera-wearer interacts with through their body or tools; (b) other objects that are relevant to the activity (e.g., supporting surfaces like kitchen top); and (c) body parts (hands and legs). Note that every time an object changes visual state (adopting the Point-of-No-Return definition from [47]), it is marked as a new object (e.g., annotators list tomato and sliced tomato as two distinct object instances).
@@
Formal task definition
Given a pair of time-synchronized egocentric and exocentric videos, as well as a query object track in one of the views, the goal is to output the corresponding mask for the same object instance in the other view for all frames where the object is visible in both views. This task is especially challenging in our dataset, since we have to handle long videos with an average length of 3 minutes, as well as very small objects with areas of only a few pixels. Importantly, note that the input to the model excludes semantic labels or names for the objects, camera pose information relating the two views, and IMU or active range sensor measurements. We do not use such information as we want to encourage the development of methods for open-world correspondence, not relying on predefined sets of objects or inputs that require non-consumer camera devices.
Metrics
We adopt the following metrics in our evaluation:
  1. 1.
  2. 2.
  3. 3.
    Contour Accuracy (CA) [117], which measures mask shape similarity after translation is applied to register the centroids of the predicted and ground-truth masks.
  4. 4.
    Visibility Accuracy (VA) [16], which evaluates the ability of the method to estimate the visibility of the object in the target view, as in practice it may often be occluded or outside the field of view. We measure this performance using balanced accuracy. Note that, in contrast to the previous metrics that compare segmentation masks at frames where the object is visible in both views, this metric is computed based on all frames with query masks.
Baselines
Spatial baseline model. This model receives as inputs an egocentric frame, the associated exocentric frame, and a query object segmentation mask in one of the views. It then outputs the mask in the other view (if the object is visible in that view). It can be thought of as a generalization of query-point correspondence approaches proposed for sparse image correspondence [65]. We implement this baseline in the form of a Transformer-based image correspondence model, XSegTx (Cross View Segmentation Transformer), which extends SegSwap [142], a method originally proposed for image co-segmentation, i.e., for segmenting common objects in a pair of images. To adapt the architecture of SegSwap for our correspondence problem, we additionally condition the model on the segmentation mask of the object of interest by feeding the query mask as a third input to the model. In particular, we first pass the egocentric frame, the exocentric frame and the query mask (as a binary mask) through a visual backbone network. We then flatten the resulting features into three sequences and pass them through the cross-image transformer with alternating self-attention and cross-attention layers. We first use the query mask features to attend to the features in the query view which are then used to cross-attend over features from the target view. This allows the model to reason over features from both views conditioned on the input mask. The resulting sequences for both views are “unflattened” and passed through a decoder to predict object segmentation masks in both views. We also pass the target view features through a classification head to classify if the query object is visible in the target view.
@@
Spatio-temporal baseline model. The spatio-temporal model receives as input the pair of ego-exo video clips as well as an object segmentation track in one of the views, and outputs segmentation masks in the other view for the frames that the object is visible in both views. It can be thought of as performing generalized tracking across views. We build our baseline model on top of XMem [24], a model originally proposed for tracking a specific target object given its segmentation mask in the first frame. In particular, our baseline model, called XView-XMem, adapts XMem to track the object across different views given ground-truth segmentation masks for one of the views in each frame. To encourage the model to learn associations of the objects between egocentric and exocentric views, we train XView-XMem to track the object in a sequence of interleaved frames of egocentric and exocentric views, i.e., each egocentric frame is followed by an exocentric frame and vice versa, as shown in Figure 27.
@@
Implementation details
For our spatial baseline model, we downsample the images to 480x480 resolution for all the views while using padding to keep the original aspect ratio of the images. For the image backbone we use the same ResNet50 [51] checkpoint as SegSwap and freeze its weights during training. Our cross-image transformer architecture also follows [142]. We use a batch size of $32$ and Adam [69] as our optimizer with a learning rate of $0.0002$ which decays to $0.0001$ after 50,000 iterations. We run all our experiments on a single Nvidia RTX A6000 GPU for 200,000 iterations.
For our spatio-temporal baseline model, we use the same visual backbone (ResNet50 [51]) and architecture as XMem [24]. Our only modification is in the information that gets inserted in the working memory at each frame. We first extract features from both ResNet and XSegTx for both both query and target frames. The corresponding features are then concatenated and projected to the original feature dimension through simple 2D convolution. We train on sequences of 8 interleaved ego and exo frames. The model is trained using AdamW as our optimizer with a learning rate of $0.00001$ for 50,000 iterations and weight decay 0.05. The batch size is 8 clip pairs. We initialize our model with the original pretrained XMem, and keep both the ResNet backbone as well as our finetuned XSegTx models frozen. Note that we do not apply any data augmentations.
@@
We use 1028 takes from the Ego-Exo dataset to train and evaluate models for this benchmark. In particular, we use the common split shared across benchmarks, with 657/838 takes for training, 156/201 takes for validation and 215/295 takes for testing in the v1/v2 version of the data release, respectively. We extract pairs of images between egocentric and exocentric views which have corresponding object masks annotated for training. This gives us a total of about 193k pairs for training from the v1 training set.
[…]
Ego XSegTx (random weights) S 50.00/50.00 0.60/0.48 0.116/0.118 0.017/0.014
Ego XSegTx S 62.63/62.40 13.88/9.11 0.154/0.177 0.239/0.161
Ego XMem (w/o finetuning) ST 42.33/48.10 13.07/11.59 0.312/0.341 0.182/0.158
Ego XView-Xmem (w/ finetuning) ST 53.28/57.31 22.14/20.67 0.176/0.192 0.325/0.311
Ego XView-Xmem (+ XSegTx) ST 54.61/52.59 22.50/7.23 0.139/0.313 0.347/0.127
Exo XSegTx (random weights) S 50.00/50.00 1.62/1.08 0.197/0.203 0.027/0.024
Exo XSegTx S 74.60/71.62 21.80/16.94 0.133/0.153 0.265/0.212
Exo XMem (w/o finetuning) ST 56.96/53.89 10.20/8.32 0.249/0.268 0.125/0.102
Exo XView-Xmem (w/ finetuning) ST 59.36/65.76 23.56/21.63 0.186/0.196 0.308/0.286
Exo XView-Xmem (+ XSegTx) ST 52.28/62.07 19.39/17.03 0.208/0.227 0.255/0.230
Table 9: Baseline evaluation on the correspondence benchmark on the v1/v2 test set, respectively. All models are trained on the v1 training set. Best results are reported in bold, second best results are underlined.
@@ Results
We break down our results across different activities in Fig. 28. We note that some activities are generally easier to model (e.g., basketball, soccer) because of limited variation in object shape and appearance whereas some activities (e.g., cooking and bike repair) are much harder to model due to larger diversity in appearance, shape and size of the objects across views. We also explicitly evaluate our baselines on their ability to predict masks for very small objects. To do so, we split our validation set based on the predicted object size in proportion to pixels in the image. We see that, all our baselines struggle on very small objects and perform increasingly well on larger object sizes. We also show some qualitative results in Fig. 30. As we can see, the spatial baseline (XSegTx) struggles to track the same object throughout the video. For example, in the bottom example, XSegTx alternates between predicting one and two object masks whereas the spatiotemporal baseline (XView-XMem) reliably tracks a single object throughout the sequence, showing the importance of exploiting temporal cues in the data.
Figure 28: Performance of both baselines per activity scenario on the v1 test set.
@@
Figure 29: Correspondence evaluated across different object sizes in the target (exo) view on the v1 test set. The object sizes range from $7e^{-6}\%$ to $11\%$ pixels in the target view.
Figure 30: Qualitative results for the different correspondence baselines.
Contributions Statement
Manolis Savva co-led the correspondence benchmark and contributed to the task definition, the annotation guidelines, the baseline design, and paper writing. Effrosyni Mavroudi co-led the correspondence benchmark and contributed to the task definition, the annotation guidelines, the baseline design, and paper writing. Lorenzo Torresani contributed to the task definition, and to editing this section. Sanjay Haresh developed the spatial baselines and contributed to data analysis, experimental results, and paper writing. Yongsen Mao developed the spatiotemporal baselines and contributed to data analysis, experimental results, and paper writing. Suyog Jain formulated the annotation pipeline, developed annotation tools and contributed to annotation guidelines and paper writing. Santhosh Ramakrishnan contributed to the annotation guidelines and the formulation of the annotation pipeline. Xitong Yang contributed to the annotation guidelines and the task definition. We would like to acknowledge Hanxiao Jiang for helpful discussions and preliminary ideas on baseline implementation. Devansh Kukreja built the render flow to generate frame-aligned videos of each camera for each take as model input.
13.A.2 Ego-Exo translation
Annotations
Translation uses the same annotations as the correspondence task.
Formal task definition
The translation benchmark focuses on generating information in the egocentric view given the exocentric view. The task is separated into two subtasks (see Fig 4): 1) Ego Track Prediction requires predicting the segmentation mask of an object in the unobserved ego frames given the object masks in an observed exo clip; 2) Ego Clip Generation entails generating RGB image values within the given masked ego view based on the exo view. For both subtasks, the input exo clip consists of 5 frames evenly sampled from a time span of 5 seconds.
Metrics
We adopt a diverse set of metrics to assess the different aspects of the generated translation. As for the task of correspondence, we use Visibility Accuracy (VA) to evaluate the ability of the method to predict the visibility of the target object in the ego view but this time given only exo frames as input. Furthermore, we adopt the following metrics defined for correspondence to gauge the performance of Ego Track Prediction: 1) Location Error (LA) 2) Intersection Over Union (IoU) and 3) Contour Accuracy (CA)  [117]. The IoU and CA are calculated after registering the centroids of the predicted mask and the ground-truth ego mask, in order to gauge mask prediction independent of location error. To evaluate Ego Clip Generation we use two popular image quality metrics (SSIM, PSNR [54]) and three perceptual metrics (DISTS [33], LPIPS [192] and CLIP similarity [124]).
Baselines
For track prediction, we implement the GAN-based method pix2pix [61] and the NeRF-based method GNT [154]. For clip generation, we employ the GAN-based method pix2pix [61] and the diffusion model DiT [116]. It is worth noting that, as discussed below, we introduce specific modifications to adapt these methods to our task requirements. All baselines utilize exo images and masks, with only the GNT model making use of the extra input of ego camera pose.
  • pix2pix-mask. We modify the generator of pix2pix to have inputs and outputs of 4 channels. Specifically, the exo frame and the exo mask are concatenated as the inputs while the 4-channel outputs are ego frame (3 channels) and ego mask (1 channel). The ego frame is supervised with the losses used in pix2pix. We use the bootstrapped cross-entropy loss [129] and the dice loss [152] for mask prediction.
  • GNT-mask. We adopt the Generalizable NeRF Transformer (GNT) [154] as another baseline leveraging the camera poses. In our adapted version, the image encoder takes a 4-channel image (exo frame and mask) as inputs to predict the ego frame and ego mask. Formally, during the training of our GNT-mask, for each point $x$ and viewing direction unit vector $d\in\mathbb{R}^{3}$, the ray transformer $f$ in GNT predicts two key attributes: RGB Color ($c$) and Object Existence Score ($e$), in which $e$ signifies the probability of an object being present at point $x$. During rendering, the volumetric radiance field encoded by the ray transformer can then be rendered into a 2D image as well as a 2D object mask.
Ego Clip Generation requires producing pixel values representing the target object in the egocentric view. To achieve this, we leverage 6 different input images for each frame: exo frame, exo mask, exo object crop, cropped exo mask, ego mask and cropped ego mask. The cropped exocentric and egocentric masks are generated by considering a bounding box to isolate the relevant portions of the exocentric and egocentric masks, respectively. The “exo object crop" refers to the RGB image obtained by cropping out the relevant region using the cropped exocentric mask. We resize these 6 images to the same size ($256\times 256$). We evaluate two baselines for this task:
  • DiT-pix. We adopt the Transformer-based diffusion model DiT [116]. We predict the ego object crop by conditioning the DiT on the 6 input images in two manners. Initially, these six images are concatenated along the channel dimension and subsequently combined with the noisy ego object crop, forming the input to DiT. Additionally, two ResNet-50 architectures encode the six images into low-dimensional features, which are then incorporated into each layer of DiT via AdaLN [118].
  • pix2pix-pix. We adopt pix2pix [61] for clip generation as well by concatenating the 6 images along the channel dimension as inputs to the pix2pix model.
All of the above-mentioned baselines perform image-to-image generation. We implement also clip-to-clip variants of these methods by taking multiple frames as inputs and predicting results for all frames jointly. For pix2pix, we achieve this by replacing the original 2D-Conv with 3D-Conv, and 2D-BatchNorm with 3D-BatchNorm. For DiT, we use space-time divided attention as in TimeSformer [13].
[…]
 pix2pix-mask No 20.7/19.6 4.1/4.0 4.6/5.4
$+$multi-frame No 23.0/18.7 8.0/9.9 8.5/13.1
GNT-mask Yes 18.5/15.5 15.7/11.8 10.1/6.6
Table 10: Evaluation of translation baselines for the subtask of ego track prediction on the v1/v2 test set, respectively. All models are trained on the v1 training set.
Results We employ the validation set for the purpose of selecting optimal checkpoints and hyper-parameters, which are subsequently evaluated on the test set.
In the context of Ego Track Prediction (Table 10), we gauge performance using Visibility accuracy (VA) to assess the ability of the model to estimate the object visibility in the ego view. Correctness is determined when two conditions are met: (1) the predicted mask is empty when the object is invisible in the ego view, and (2) the predicted mask is non-empty when the object is visible in the ego view. Both pix2pix-mask and GNT-mask fail to perform well in estimating the object visibility, achieving Visibility accuracy about the same as the 50% performance of random guess (55.6/50.0% for GNT-mask and 47.6/56.2% for pix2pix-mask, on the v1/v2 test set, respectively). However, the ResNet-50 trained exclusively to attend to this binary classification achieves a VA of 80.6/77.5% on the v1/v2 test set, respectively. We assess mask quality by considering distance (Location Error) and similarity metrics (IoU and Contour Accuracy) between predicted and ground-truth masks after registration. The 3D-aware NeRF-based baseline, GNT-mask, outperforms the implicit baseline, pix2pix-mask, overall. However, it does so by exploiting the ego camera pose as additional input. It is noteworthy that both baselines perform poorly on this task, likely due to the inherent challenges in correctly predicting the location and shape of the target object in the ego view, probably due to the fact that it often has diminutive size in the exo view.
 pix2pix-pix 0.51/0.41 16.2/17.3 0.37/0.36 0.60/0.49 72.1/77.4
DiT-pix 0.56/0.64 16.0/19.0 0.31/0.27 0.47/0.38 83.6/83.5
Table 11: Evaluation of translation baselines for the subtask of ego clip generation on the v1/v2 test set, respectively. All models are trained on the v1 training set.
In the case of Ego Clip Generation (Table 11), the Diffusion model DiT-pix demonstrates superior performance across all metrics compared to the GAN-based pix2pix-pix. Qualitative results (Figure 31) illustrate that DiT-pix can generate highly photorealistic images, aligning closely with the ground-truth in most instances. However, there are occasional cases (the last 2 rows) where the shape of the object is accurately generated, but the texture deviates slightly.
We further verify the importance of each input in Figure 32. Without exo object crop as input, the model fails to correctly infer the color and texture of the target object in the ego view. This result is expected as the source objects often represent a very small region of the entire exo frame. Additionally, without the ego crop mask as input, the model predicts the orientation of the object incorrectly. These observations highlight the importance of the cropped inputs.
We can observe in Table 10 that multi-frame (i.e., clip-to-clip) prediction does not provide a quantitative advantage over frame-to-frame prediction. Yet, we noticed that the multi-frame variant often yields generations that are more consistent across frames, even for frames where the exo view is heavily occluded, as can be seen in Figure 33. This is reasonable as a clip-level model can more effectively learn about the target object from multiple frames and fill-in information that is missing in individual exo frames.

### v4

@@ Appendix D Language Descriptions
As introduced in the main paper, Ego-Exo4D provides three forms of parallel text corpora for the video: expert commentary, narrate and act, and atomic action descriptions. Table 17 shows examples from different scenarios highlighting their distinctions in style and point of view. Figure 33 shows word clouds per scenario and annotation type highlighting the differences in vocabulary and word frequency.
D.1 Expert Commentary Tool
To collect expert commentary, we developed a web-based tool, which is open sourced as part of the Ego-Exo4D dataset and benchmark suite. See Figure 35. Known as the Narrator, this application supports video playback for Ego-Exo4D skills demonstrations, records time-stamped verbal commentary, and allows exporting and viewing commented videos. As a web-based platform, the Narrator can be simply accessed through a browser, with minimal set-up and less restrictive system requirements compared to tools requiring local installation. These attributes made it efficient to onboard and manage our geographically distributed experts. We acknowledge the EPIC Narrator (Damen et al., 2018) as the open-sourced inspiration and source code for this initiative.
@@ D.1 Expert Commentary Tool
D.2 Atomic Action Description Statistics
See Table 18 for atomic action descriptions summary statistics.
@@ D.2 Atomic Action Description Statistics
Table 18: Atomic action descriptions per-domain statistics.
Appendix E Benchmarks: annotations and baselines
We employed a unified dataset split for EgoExo4D applicable for all benchmarks. Splits were defined at the “take” level, i.e. each take was allocated to either the training, validation, or testing set. Any derivatives from individual takes (e.g. segment clips) simply inherit the split assignment from the original take.
Additionally, the splits were stratified according to activity types and proficiency scores. This was done to ensure a balanced distribution of data across the training, validation, and testing sets, with proportional representation of activities and proficiency levels. To prevent data leakage, participants and all their associated artifacts were assigned exclusively to one of the splits. This constraint was particularly important because several participants contributed multiple takes, and it was crucial to avoid sharing their data between the splits. In the end we divided a total of 5045 takes into 3082 training, 842 validation and 1121 testing takes.
E.1 Relation
@@
  • Stage 1: Egocentric video annotation. Annotator watches the egocentric video and is also shown (a) text and (b) a bounding box for one of the objects annotated in the previous stage. Annotator then marks a segmentation mask for that object in all the video frames where the object is visible. Segment Anything (Kirillov et al., 2023) is leveraged to generate segmentation masks efficiently using only point clicks.
@@
What are the objects of interest? We focus on objects that are active at some point during the execution of the activity. These objects are not only interesting because they are essential to the activity, but they are also challenging to track, since they are moving/changing state. In particular, our annotation guidelines requested annotators to list (a) objects that the camera-wearer interacts with through their body or tools; (b) other objects that are relevant to the activity (e.g., supporting surfaces like kitchen top); and (c) body parts (hands and legs). Note that every time an object changes visual state (adopting the Point-of-No-Return definition from (Grauman et al., 2022)), it is marked as a new object (e.g., annotators list tomato and sliced tomato as two distinct object instances).
@@
E.1.1 Ego-exo correspondence
Correspondence Baseline Implementation Details
Spatial baseline model. To adapt the architecture of SegSwap (Shen et al., 2022) for our correspondence problem, we additionally condition the model on the segmentation mask of the object of interest by feeding the query mask as a third input to the model. In particular, we first pass the egocentric frame, the exocentric frame and the query mask (as a binary mask) through a visual backbone network. We then flatten the resulting features into three sequences and pass them through the cross-image transformer with alternating self-attention and cross-attention layers. We first use the query mask features to attend to the features in the query view which are then used to cross-attend over features from the target view. This allows the model to reason over features from both views conditioned on the input mask. The resulting sequences for both views are “unflattened” and passed through a decoder to predict object segmentation masks in both views. We also pass the target view features through a classification head to classify if the query object is visible in the target view.
@@ E.1.1 Ego-exo correspondence
Spatio-temporal baseline model. To encourage the model to learn associations of the objects between egocentric and exocentric views, we train XView-XMem to track the object in a sequence of interleaved frames of egocentric and exocentric views, i.e., each egocentric frame is followed by an exocentric frame and vice versa, as shown in Figure 37.
@@
For our spatial baseline model, we downsample the images to 480x480 resolution for all the views while using padding to keep the original aspect ratio of the images. For the image backbone we use the same ResNet50 (He et al., 2016) checkpoint as SegSwap and freeze its weights during training. Our cross-image transformer architecture also follows (Shen et al., 2022). We use a batch size of $32$ and Adam (Kingma and Ba, 2014) as our optimizer with a learning rate of $0.0002$ which decays to $0.0001$ after 50,000 iterations. We run all our experiments on a single Nvidia RTX A6000 GPU for 200,000 iterations.
For our spatio-temporal baseline model, we use the same visual backbone (ResNet50 (He et al., 2016)) and architecture as XMem (Cheng and Schwing, 2022). Our only modification is in the information that gets inserted in the working memory at each frame. We first extract features from both ResNet and XSegTx for both both query and target frames. The corresponding features are then concatenated and projected to the original feature dimension through simple 2D convolution. We train on sequences of 8 interleaved ego and exo frames. The model is trained using AdamW as our optimizer with a learning rate of $0.00001$ for 50,000 iterations and weight decay 0.05. The batch size is 8 clip pairs. We initialize our model with the original pretrained XMem, and keep both the ResNet backbone as well as our finetuned XSegTx models frozen. Note that we do not apply any data augmentations.
@@
We use 1028 takes from the Ego-Exo dataset to train and evaluate models for this benchmark. In particular, we use the common split shared across benchmarks, with 838 takes for training, 201 takes for validation and 295 takes for testing. We extract pairs of images between egocentric and exocentric views which have corresponding object masks annotated for training. This gives us a total of about 193k pairs for training.
@@ Results
We break down our results across different activities in Fig. 38. We note that some activities are generally easier to model (e.g., basketball, soccer) because of limited variation in object shape and appearance whereas some activities (e.g., cooking and bike repair) are much harder to model due to larger diversity in appearance, shape and size of the objects across views. We also explicitly evaluate our baselines on their ability to predict masks for very small objects. To do so, we split our validation set based on the predicted object size in proportion to pixels in the image. We see that, all our baselines struggle on very small objects and perform increasingly well on larger object sizes.
Figure 38: Performance of both baselines per activity scenario.
@@
Figure 39: Correspondence evaluated across different object sizes in the target (exo) view. The object sizes range from $7e^{-6}\%$ to $11\%$ pixels in the target view.
E.1.2 Ego-exo translation